OpenStudy (anonymous):

, Find dy/dx if x^2+y^2=2xy... Please and Thank you :)

OpenStudy (valpey):

\[(x-y)^2 = 0\] \[y=x\] \[\frac{dy}{dx} = 1\]

OpenStudy (anonymous):

2x+2yy'=2y+2xy' x+yy'=y+xy' x-y=(x-y)y' y'=1

OpenStudy (valpey):

Gotta love them trick questions.

OpenStudy (anonymous):

Yup.... I know I absolutely adore them -_-

OpenStudy (anonymous):

Thank you guys :D